Grade 11Mechanics

a ball is thrown vertically upwards with a velocity 'u' from the balloon descending with a constant velocity 'v'.. the ball will pass by the balloon after what time?

mohammad inaam ul haque
9 Years agoGrade 11

1 Answer

Khimraj
9 Years ago
Write the equation of displacement for bothFor ball s1 =- ut+1/2gt²For ballon s2 = vtNow according to question s1 = s2(v+u)t = 1/2(gt²)t = 2(u+v)/gPlease if u like answer then approve answer